An athlete buys three T-shirts, five sweat-bands and two pairs of running shoes from a sports outlet.

A T-shirt costs twice as much as a sweat-band, while a pair of running shoes costs $ 300 more than a

sweat-band. The athlete hands the cashier $ 1380

Determine the cost of a T-shirt, a sweat-band and a pair of running shoes.


1
Expert's answer
2021-11-15T16:11:56-0500

Three T-shirts cost (the cost of each is x):


"C_1=3x"

Five sweat-bands cost (the cost of each is y):


"C_2=5y"

Two pairs of running shoes (the cost of a pair is z):


"C_3=2z"



Now, translate to math language the following phrases.

A T-shirt costs twice as much as a sweat-band (equation 1):


"x=2y."

A pair of running shoes costs $ 300 more than a sweat-band (equation 2):


"z=300+y."

The athlete hands the cashier $ 1380:


"C_1+C_2+C_3=1380,\\
3x+5y+2z=1380."

Substitute equation 1 and equation 2 into the last equation:


"3(2y)+5y+2(300+y)=1380,\\
6y+5y+600+2y=1380,\\
13y+600=1380,\\
13y=780,\\
y=60.\\\space\\
x=2y=2·60=120,\\
z=300+y=300+60=360."